# Refined Bitcoin Security-Latency Under Network Delay

### v2

@@ 4.2 Confirmation Interval
Similar to the case of the lower bound we find the number of blocks mined before each honest jumper is published. Keep in mind that, in addition to delaying the publication of jumpers by $\Delta$, all honest blocks mined during this delay are rigged. Hence, all blocks but jumpers are adversarial in this model.
@@ 4.3 Post-Confirmation Race
Clearly, if $\bar{L}+\bar{S}\geq k$, then, the adversary wins the race, otherwise, the deficit of the adversary is $\bar{D}=k-\bar{L}-\bar{S}$, which it has to make up during the post-confirmation race. Here, we go back to the simplified $(A,H)$ random walk model and consider two arrivals at a time as we did for the lead. It is straightforward to show that a random walk which starts at the origin and moves two steps at a time, left or right with $\mathbb{P}(\bar{E_{1}})=\bar{\alpha}^{2}$ and $\mathbb{P}(\bar{E_{3}})=\bar{\beta}^{2}$, respectively, can reach the point $2m$ with probability
@@ Appendix A Appendix
Parameters Values
$a_{i}$ $\sum_{j\geq i}\alpha_{j}+\beta*\mathbbm{1}_{i\leq 2}$
$b_{i}$ $\sum_{j\geq i}\alpha_{j}+\beta*\mathbbm{1}_{i\leq 1}$
$\bar{a}_{i}$ $\sum_{j\geq i}\bar{\alpha}_{j}+\beta*\mathbbm{1}_{i\leq 2}$
$\bar{b}_{i}$ $\sum_{j\geq i}\bar{\alpha}_{j}+\beta*\mathbbm{1}_{i\leq 1}$
Table 2: Ramaswami parameters.
The steady state of matrix $P$ can be found by Ramaswami’s formula [8], as $P$ is a stochastic matrix of M/G/1 type.
@@
Proof:  Note that, finding $a_{i}$ and $b_{i}$ is a straightforward application of Ramaswami’s formula and they are provided in Table 2. Here, we only prove $\pi_{0}=\frac{1-\beta(2+\alpha\lambda\Delta)}{\alpha}$.
Summing (71) for $i\geq 1$, we get
$(1-a_{1})\sum_{i\geq 1}\pi_{i}$ $=\sum_{i\geq 1}\left(\pi_{0}b_{i}+\sum_{j=1}^{i-1}\pi_{j}a_{i+1-j}\right)$ (72)
$=\pi_{0}\sum_{i\geq 1}b_{i}+\sum_{i\geq 2}\sum_{j=1}^{i-1}\pi_{j}a_{i+1-j}$ (73)
$=-\pi_{0}\beta+\pi_{0}\sum_{i\geq 1}a_{i}+\sum_{j\geq 1}\sum_{i>j}\pi_{j}a_{i+1-j}$ (74)
$=-\pi_{0}\beta+\pi_{0}\sum_{i\geq 1}a_{i}+\sum_{j\geq 1}\pi_{j}\sum_{i\geq 2}a_{i}$ (75)
Substituting $\sum_{i\geq 1}\pi_{i}=1-\pi_{0}$ gives
$(1-a_{1})(1-\pi_{0})=-\pi_{0}\beta+\pi_{0}\sum_{i\geq 1}a_{i}+(1-\pi_{0})\sum_{i\geq 2}a_{i}$ (76)
and moving terms, we have
$\alpha\pi_{0}$ $=1-\sum_{i\geq 1}a_{i}$ (78)
$=1-\sum_{i\geq 1}\sum_{j\geq i}\alpha_{j}-2\beta$ (79)
$=1-\sum_{j\geq 1}\sum_{i=1}^{j}\alpha_{j}-2\beta$ (80)
$=1-\sum_{j\geq 1}j\alpha_{j}-2\beta$ (81)
$=1-\alpha\beta\lambda\Delta-2\beta$ (82)
gives the desired result. $\blacksquare$
@@
Corollary 4
Steady state of $\bar{P}$ can be found recursively using
where $\pi_{0}=\frac{1-2\beta-\alpha\lambda\Delta}{\alpha}$.
Proof:  The proof follows the same steps as in Lemma 5, where we replace $a_{i},b_{i}$ and $\alpha_{i}$ with $\bar{a}_{i},\bar{b}_{i}$ and $\bar{\alpha}_{i}$. After (81), we end up with $\alpha\lambda\Delta$ instead of $\alpha\beta\lambda\Delta$, since we consider all arrivals for each $\Delta$ delay interval instead of just adversarial arrivals. $\blacksquare$
@@
Further, we cannot claim that the lead of $\bar{P^{\prime}}$ upper bounds the lead of $\bar{P}$, i.e., $P(\bar{L^{\prime}}<i)\leq P(\bar{L}<i)$ is not always true. When we derived $P^{\prime}$ from $P$, we truncated number of adversarial arrivals, which makes sure $L^{\prime}\leq L$ for all arrival paths (i.e., $L^{\prime}\leq L$ almost surely). However, the same is not true for $\bar{P^{\prime}}$ as we change the dynamics of the walk. For example, consider first four arrivals after genesis block with $t_{1}>\Delta$, $t_{2}<t_{1}+\Delta$ both honest and $t_{3}>t_{2}+\Delta$ and $t_{4}>t_{3}$ both adversarial. The lead model presented by $\bar{P}$ follows the path $0\rightarrow 0\rightarrow 1\rightarrow 2\rightarrow 3$. However, the lead presented by $\bar{P^{\prime}}$ for the same arrival sequence follows the path $0\rightarrow 1\rightarrow 0\rightarrow 1\rightarrow 2$. Hence, $\bar{L}<\bar{L^{\prime}}$ does not necessarily hold. As a result, the bounds of Theorem 2 obtained by $\bar{P}$ and $\bar{P^{\prime}}$ differ for different values of $k$, and one can compare them to pick the better bound for each $k$.
In our figures, we use the steady states of $P^{\prime}$ and $\bar{P^{\prime}}$ consistently. We note that the bounds of the two lower bounds visually overlap with each other, and so do the two upper bounds, hence we refrain from comparing them with plots. As an example, for Bitcoin $6$-block confirmation rule with $\alpha=90$, using $\bar{P^{\prime}}$ results in safety violation upper bound probability of $0.00173$ whereas $\bar{P}$ gives $0.00177$. For the lower bound, the values are even closer, with $P^{\prime}$ and $P$ resulting in $0.00112412$ and $0.00112415$, respectively.

### v3

@@ 4.2 Confirmation Interval
As the transaction enters the mempool at $\tau$, it will be included in the next honest block to be mined. However, if $\tau$ is in the $\Delta$ delay immediately following a jumper that arrived before $\tau$, the honest blocks mined during this $\Delta$ delay are rigged. Hence, since the $tx$ will be included in the first jumper block after $\tau$, without loss of generality, we assume again that $tx$ enters the mempool at a time $\tau$, where the first honest block to be mined contains the $tx$ and is a jumper. Further, this implies that the steady-state distribution of $\bar{P}$ is an upper bound for the adversarial lead. Similar to the case of the lower bound we find the number of blocks mined before each honest jumper is published. Keep in mind that, in addition to delaying the publication of jumpers by $\Delta$, all honest blocks mined during this delay are rigged. Hence, all blocks but jumpers are adversarial in this model.
@@ 4.3 Post-Confirmation Race
Clearly, if $\bar{L}+\bar{S}\geq k$, then, the adversary wins the race, otherwise, the deficit of the adversary is $\bar{D}=k-\bar{L}-\bar{S}$, which it has to make up during the post-confirmation race.
Here, we consider two arrivals at a time which brings improvement to the random walk considered in [5]. In the rigged model of [5], an arrival is i.i.d. Bernoulli, and coin toss lands honest w.p. $\alpha e^{-\lambda\Delta}$. Let $(H,A)$ denote the growth of honest and adversarial chains after two arrivals. If we consider two consecutive i.i.d. coin tosses separately and sum their results, then, 1. $\bar{E}_{1}^{\prime}$denoting $(2,0)$ happens w.p. $(\alpha e^{-\lambda\Delta})^{2}$; 2. $\bar{E}_{2}^{\prime}$denoting $(1,1)$ happens w.p. $2(\alpha e^{-\lambda\Delta})(1-\alpha e^{-\lambda\Delta})$; and 3. $\bar{E}_{3}^{\prime}$denoting $(0,2)$ happens w.p. $(1-\alpha e^{-\lambda\Delta})^{2}$. In our analysis $\bar{E}_{1}^{\prime}$ stays the same but we decrease the probability of $\bar{E}_{3}^{\prime}$ by considering two consecutive coin tosses together instead of considering them separately.
Without loss of generality, assume that we start from time zero and consider groups of two arrivals in the rigged model. Note that the genesis block is the zeroth jumper that arrives at time zero. Here, $\bar{E}_{1}$, i.e., $(H,A)=(2,0)$ has the same probability as $\bar{E}_{1}^{\prime}$. For $\bar{E}_{2}$ to happen, i.e., $(H,A)=(1,1)$, there are $3$ cases:
  1. 1.
  2. 2.
  3. 3.
After the second arrival at $t_{2}$, by the memorylessness property of exponential arrivals, same arguments will hold for the next group of two arrivals and so forth. Hence, each group is identically distributed and independent from each other. Note that, $\mathbb{P}(\bar{E}_{2})=\rho\geq\mathbb{P}(\bar{E}_{2}^{\prime})$, hence the improvement.
It is straightforward to show that a random walk which starts at the origin and moves two steps at a time, left or right with $\mathbb{P}(\bar{E_{1}})=\bar{\alpha}^{2}$ and $\mathbb{P}(\bar{E_{3}})=\bar{\beta}^{2}$, respectively, can reach the point $2m$ with probability
@@ Appendix A Appendix
Proof:  The proof follows the same steps as in Lemma 4, where we replace $\bar{a}_{i}$, $\bar{b}_{i}$ and $\bar{\alpha}_{i}$ with $a_{i}$, $b_{i}$ and $\alpha_{i}$. After (45), we end up with $\alpha\beta\lambda\Delta$ instead of $\alpha\lambda\Delta$, since we consider just adversarial arrivals for each $\Delta$ delay interval instead of all arrivals. $\blacksquare$
@@
Further, with the random walk presented in $P$, the random walk escapes to infinity when $1\leq\beta(2+\alpha\lambda\Delta)$ which is the same condition as $\beta\geq\frac{1-\beta}{1+(1-\beta)\lambda\Delta}$, i.e., the converse of the ultimate fault tolerance condition. This is not surprising, since $P$ presents the $\Delta$ delay strategy, where the honest chain grows at the slowest possible rate, i.e., $\frac{1}{\Delta+1/\alpha\lambda}$, and is in a race with adversarial chain that has the rate $\beta\lambda$.
In our figures, we use the steady states of $P^{\prime}$ and $\bar{P}$ consistently. We note that the bounds of the two lower bounds visually overlap with each other, hence we refrain from comparing them with plots. As an example, for Bitcoin $6$-block confirmation rule with $\alpha=0.9$, using $P^{\prime}$ results in safety violation lower bound probability of $0.00112412$ whereas $P$ gives $0.00112415$.